Suppose the following equations characterize the supply and demand in the labor
market model:
Labor Supply: LS = 2W+30
Labour Demand: LD = 60 - w(1+T)
W = WAGE/ T = PAYROLL TAX.

Suppose T= 0.2, solve for the equilibrium values of wage and labour supply.
1
Expert's answer
2014-09-24T11:12:53-0400
Labor Supply: LS = 2W+30
Labour Demand: LD = 60 - w(1+T)
W = w/ T = PAYROLL TAX.
T= 0.2,
We can find equilibrium values of wage and labour supply in the point, where LS = LD
2W + 30 = 60 - w(1+T)
2w/0.2 + 30 = 60 - w(1+T)
10w + 30 = 60 - w + 0.2w
10.8w = 30
w = 30/10.8 = $2.78
LS = 2*2.78/0.2 + 30 = 57.8 = 58 workers.